# Axi-majoron for almost everything

### v2

@@ 3 Inflation along $\Phi$ direction
Therefore, for a sufficiently small $\xi_{\chi}$ with $\xi_{\phi}\gtrsim\mathcal{O}(1)$, it is possible to have $\sqrt{\xi_{\phi}}\phi_{\rm ini}\gg M_{\rm P}$ while Planck scale is nearly fixed, and we assume this is the case. In this work, we will be interested in $\xi_{\phi}=\mathcal{O}(1-10^{2})$, concerning various cosmological aspects of this model. Hence, the unitarity-cutoff may be slightly lower than the Planck scale, but it is well above any energy and mass scales involved in the field dynamics. Also, even though the field excursion during inflation spans over the Planck scale as will be discussed subsequently, it may present no problem as long as scale-invariance is preserved.
In the large field region around or over the Planck scale, the kinetic terms of the scalar fields $\phi_{r}$ and $\phi_{i}$ in Einstein frame are neither canonical nor diagonal. In order to analyze the dynamics of fields in such a curved field-space, we resort to the conventional method for multi-field inflation in a curved field-space [38, 39]. From now on we work in the basis of $\phi^{a=1,2}\equiv(\phi_{r},\phi_{i})^{T}$ unless otherwise stated. We define the Levi-Civita connection, covariant field-derivative $\nabla_{a}$, and covariant time-derive in the field space as
@@ 3.1 Dynamics of the homogeneous mode
One can solve numerically these equations of motion of fields straightforwardly. However, it is instructive to see a limiting case which can be proven simpler to understand. For that, one can take a look at the case of $\alpha\lll 1$ for whose(?) approximate analytic expressions can be obtained as follows. The frame function in $\left(\phi,\theta\right)$-basis is approximated as
@@
$K_{22}$$\approx$ $1+\frac{\xi_{\phi}\left[1+6\xi_{\phi}\left(\alpha\cos(2\theta)\right)^{2}\right]\phi^{2}}{M_{\rm P}^{2}}\ ,$
$K_{12}$$\approx$ $6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\alpha\cos\left(2\theta\right)\ll 1\ .$ (3.13)
@@
$H_{I}\simeq\frac{\lambda_{\Phi}^{1/2}M_{\rm P}}{2\sqrt{3}\xi_{\phi}}$ (3.20)
@@
From Eq. (3.22), the mass-squard along $\varphi_{2}$-direction is found to be
@@
$\alpha_{\rm c}\equiv 1/(24\xi_{\phi}),$ (3.25)
the mass along $\varphi_{2}$ direction (i.e., angular direction in the complex field space of $\Phi$) is comparable to the expansion rate given in Eq. (3.20). This implies that, in the large field region where the potential along the radial direction (i.e., along $\varphi_{1}$ or $\phi$) becomes quite flat, if $\alpha\gtrsim\alpha_{\rm c}$, the dynamics of field would be initially along $\varphi_{2}$ toward the vicinity of $\varphi_{2}\sim 0$ and afterward moves nearly along $\varphi_{1}$ toward the origin, as will be shown by numerical analysis. On the other hand, if $\alpha\lll\alpha_{\rm c}$, the trajectory will be nearly radial towards the origin. That is, the trajectory of the field would look like one of the following three cases in $(\phi_{r},\phi_{i})$ space:
@@
Figure 2: Trajectories of the field dynamics for $\alpha=10^{-4},10^{-3},10^{-2},10^{-1}$ from left to right with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/16,\pi/4,7\pi/16$(from left to right).
@@
As it can be seen from the dynamics of the inflaton with $\theta_{\rm ini}\neq 0$, a salient and interesting feature of our hybrid axi-majoron inflation is that the PQ field $\Phi$ can have angular motion even after inflation, if $\alpha$ is neither too large nor too small. Fig. 3 shows the time-dependence of $\theta$ and its time-derivative.
Figure 3: The time dependence of the phase $\theta$ and its time-derivative for $\alpha=10^{-3}$ with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$. $H_{\rm ref}\equiv\sqrt{U(\phi_{\rm ini},\theta_{\rm ini})}/\sqrt{3}M_{\rm P}\approx H_{I}$. The vertical dotted line indicates the end of inflation.
The angular momentum in a co-moving volume is conserved after inflation, since the symmetry-breaking effect becomes negligible as $\phi$ becomes smaller than thePlanck scale. This corresponds to the asymmetry of $U(1)_{\rm PQ}$-charged particle $\Phi$ as in the case of Affleck-Dine mechanism [29]  [^13] In Ref. [41], generation of Peccei-Quinn number asymmetry and subsequent baryogenesis was discussed, but in a different realization.. The asymmetry can be transferred to the visible sector, as the baryon number asymmetry, via the usual way of Affleck-Dine mechanism and/or the mechanism of spontaneous baryogenesis [42]. More details will be discussed in Section 4.
@@ 3.2 Perturbations
For the estimations of perturbations [^14] There are perturbations of $\chi$ field too. However, for $\xi_{\chi}\ll\xi_{r,i}$ the background dynamics is nearly along $\Phi$ with barely changing $\chi$. In addition, in a simplified two field case with negligible kinetic mixing which is the case of $\xi_{\chi}\lll 1$, one can see that the growing rate of iso-curvature perturbation is negative (i.e., suppression instead of growth) in the slow-roll and slow-turn limit [43]. Hence, we can ignore the perturbations of $\chi$ in our discussion. , we take the Newtonian gauge in an isotropic universe for the metric perturbations, i.e., the perturbed metric is given by
@@
$\epsilon$$\equiv$ $-\frac{\dot{H}}{H^{2}}=\frac{\dot{\phi}_{I}^{2}}{2H^{2}}$ (3.30)
@@
¨$\ddot{\mathcal{R}}+\left(3+2\epsilon-2\eta_{T}\right)H\mathcal{R}+\frac{k^{2}}{a^{2}}\mathcal{R}$ $=$ $-2\dot{\theta}_{T}\frac{H}{\dot{\phi}_{I}}\left[\dot{Q}_{N}+\left(3-\eta_{T}-2\epsilon+\frac{\dot{\eta}_{N}}{H\eta_{N}}\right)HQ_{N}\right]$ (3.37)
@@
For numerical analysis, we took $\xi_{\phi}=1$, $\alpha=10^{-3}$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$ as an illustrative set of parameters.
@@
Figure 4: Left: The turning rate, $\dot{\theta}_{T}(t)$, for $t\leq t_{e}$. Right: The effective mass-squared $m_{\rm eff}^{2}$ associated with iso-curvature perturbations for the same parameter set as the one used in Fig. 3. The vertical dotted line corresponds to $52$ $e$-folds to the end of inflation at $t=t_{e}\simeq 352H_{\rm ref}^{-1}$. Roughly, it corresponds to the time when the co-moving scale associated with the pivot scale of the Planck mission exits the horizon during inflation.
@@
In Fig. 4, $\dot{\theta}_{T}$ and $m_{\rm eff}^{2}$ are depicted. From the left panel of the figure, one can see that, $|\dot{\theta}_{T}|=\mathcal{O}(10^{-3})\times H_{I}$, i.e., the turning rate of the inflaton trajectory is quite small. Also, the right panel shows that $m_{\rm eff}^{2}$ is not negligible and positive at least for the last about 60 $e$-folds of inflation, leading to a suppression of iso-curvature perturbations. As a result, the source terms in both of Eqs. (3.37) and (3.38) (i.e., the right-hand sides of the equations) are negligible, and there is no sizable impact of a two-dimensional curvilinear trajectory of the inflaton on the evolution of adiabatic and isocurvature perturbations associated with cosmological scales of the present universe. Even though the inflaton trajectory involves axi-majoron direction which has anomalous couplings to non-Abelian gauge fields of the SM, the perturbations of gauge fields are not enhanced because the velocity of axi-majoron during inflation is maintained to be smaller than the expansion rate by at least a couple of orders of magnitude during the whole relevant $e$-folds, as shown in Fig. 3. Non-Gaussianities are not enhanced [48] either. In Fig. 5, we depict adiabatic (left) and iso-curvature perturbations (right) for comoving scales exiting horizon at $N\approx 37$(dashed red) and $52$(solid blue) as examples.
Figure 5: Adiabatic and iso-curvature perturbations of the mode $k_{N}=a_{N}H_{I}$ for $N=37$(dashed red) and $52$(solid blue) with the same parameter set as the one used in Fig. 3.
@@
$P_{R}$ $=$ $\frac{1}{8\pi^{2}\epsilon}\times\frac{H^{2}}{M_{\rm P}^{2}}\ ,$ (3.42)
@@
For the potential given in Eq. (3.19) with $\alpha\ll 1$, the relations between slow-roll parameters and $e$-folds $N_{e}$ in the SM Higgs-inflation [40] are expected still to hold at least approximately. The latter is given by [40]
$N_{e}(\phi)\approx\frac{3}{4}\frac{\phi^{2}-\phi_{e}^{2}}{M_{\rm P}^{2}/\xi_{\phi}}$ (3.45)
as the $e$-folds from the end of inflation at $\phi_{e}\approx\left(4/3\right)^{1/4}M_{\rm P}/\sqrt{\xi_{\phi}}$. Slow-roll parameters are then given by
$\epsilon$ $\simeq$ $\frac{3}{4}N_{e}^{-2}\ ,$ (3.46)
$\eta_{T}$ $\simeq$ $N_{e}^{-1}\ ,$ (3.47)
@@
$P_{R}$ $\simeq$ $\frac{\lambda_{\phi}N_{e}^{2}}{72\pi^{2}\xi_{\phi}^{2}}\ ,$ (3.48)
$n_{s}$ $\simeq$ $1-2N_{e}^{-1}\ ,$ (3.49)
$r_{T}$ $\simeq$ $12N_{e}^{-2}\ .$ (3.50)
Density perturbations of the universe extracted from various observations including CMB observations lead to $P_{R}\simeq 2.1(\pm 0.03)\times 10^{-9}\ (1\sigma\ {\rm CL})$, $n_{s}=0.9658\pm 0.004\ (1\sigma\ {\rm CL})$, and $r_{T}<0.068\ (2\sigma\ {\rm CL})$ for the pivot scale, $k_{*}=0.05/{\rm Mpc}$ [51]. Hence, Eq. (3.48) constrains $\lambda_{\phi}$ and $\xi_{\phi}$ to satisfy a relation,
$\lambda_{\phi}\sim 6\times 10^{-10}\xi_{\phi}^{2}\left(\frac{52}{N_{e}}\right)^{2}\ .$ (3.51)
where we used $N_{e}=52$ as a representative value associated with $k_{*}$. In Table 2, we show a result of numerical estimations for some sets of parameters set as
$\lambda_{\phi}\equiv\lambda_{\phi}^{\rm ref}\times\xi_{\phi}^{2},\ \alpha=\frac{10^{-3}}{\xi_{\phi}},\ \phi_{\rm ini}=\frac{20M_{\rm P}}{\sqrt{\xi_{\phi}}},\ \theta_{\rm ini}=\frac{\pi}{4}\ ,$ (3.52)
with $\lambda_{\phi}^{\rm ref}\equiv 6\times 10^{-10}$ and several different values of $\xi_{\phi}$ but $N_{e}=52$ fixed.
$\xi_{\phi}$ $P_{R}\times 10^{9}$ $n_{s}$ $r_{T}\times 10^{3}$
@@
The result is in a good agreement with approximations in Eq. (3.48) – (3.50), and can match well with observations by adjusting $\lambda_{\phi}$. The sizable difference between the cases of $\xi_{\phi}=1$ and the other much larger values may be understood from the dependence of $U$ on $\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ in Eq. (3.19). For convenience, in the subsequent argument we use $\lambda_{\phi}$ parameterized as in Eq. (3.52).
@@
$\xi_{\phi},\lambda_{\phi},\alpha,\theta_{\rm ini},\phi_{\rm ini}\ .$ (3.53)
Once $\xi_{\phi}$ as a free parameter is chosen such that $\xi_{\phi}\gtrsim\mathcal{O}(1)$ [^15] As long as $\lambda_{\phi}\lll 1$ with $\xi_{\phi}$ given by the relation Eq. (3.51), the expansion rate during inflation and the maximal energy scale right after inflation are always well below the scale of the unitarity cutoff, $\Lambda_{\rm cut}\sim M_{\rm P}/\sqrt{\xi_{\phi}}$ [52, 53, 54]. , $\lambda_{\phi}$ is nearly fixed as long as $\alpha$ is sufficiently smaller than unity. Dependence of observables on the other parameters for given $\xi_{\phi}$ and $\lambda_{\phi}$ is shown in Fig. 6.
Figure 6: $P_{R}$(left) and $n_{s}$(right) as functions of $\theta_{\rm ini}$ for various values of $\alpha$ with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}/M_{\rm P}=10({\rm top}),20({\rm bottom})$.
It turns out that inflationary observables are barely affected by $\theta_{\rm ini}$ for $\sqrt{\xi_{\phi}}\phi_{\rm ini}\gtrsim 10M_{\rm P}$ as long as $\alpha\xi_{\phi}\lesssim\mathcal{O}(10^{-2})$ except the case of $\theta_{\rm ini}$ close to $\pi/2$ with $\sqrt{\xi_{\phi}}\phi_{\rm ini}\sim 10M_{\rm P}$ when $\alpha\xi_{\phi}\sim 10^{-2}$. Note that, if angular dynamics is not strong, for a given $e$-folds associated with the pivot scale of observations, the spectral index is nearly fixed irrespective of initial position of the field dynamics. The amplitude of power spectrum has only minor dependence on the initial position, and the dependence can be compensated by adjusting $\lambda_{\phi}$ without affecting the spectral index. This is because, while the former is a function of $e$-folds only, the latter is a function of $e$-folds and energy density of inflation. On the other hand, if $\alpha\xi_{\phi}\gtrsim\mathcal{O}(10^{-1})$, the observables varies a lot, depending on $\theta_{\rm ini}$, but $\phi_{\rm ini}$-dependence becomes weak if it is sufficiently larger than the minimal value to accommodate about 60 $e$-folds.

### v3

@@ 3 Inflation along $\Phi$ direction
Therefore, for a sufficiently small $\xi_{\chi}$ with $\xi_{\phi}\gtrsim\mathcal{O}(1)$, it is possible to have $\sqrt{\xi_{\phi}}\phi_{\rm ini}\gg M_{\rm P}$ while Planck scale is nearly fixed, and we assume this is the case.
In this work, we will be interested in $\xi_{\phi,0}=\mathcal{O}(1-10^{2})$, concerning various cosmological aspects of this model. Hence, if $c_{\phi}\chi_{0}\ll M_{\rm P}$ in Eq. (2.7), the unitarity-cutoff might be slightly lower than the Planck scale (e.g., $\Lambda_{\rm cut}\sim M_{\rm P}/\sqrt{\xi_{\phi,0}}$ [39, 40, 41]), but it is well above any energy and mass scales involved in the field dynamics. Also, even though the field excursion during inflation spans over the Planck scale as will be discussed subsequently, it may present no problem as long as scale-invariance is preserved.
In the large field region around or over the Planck scale, the kinetic terms of the scalar fields $\phi_{r}$ and $\phi_{i}$ in Einstein frame are neither canonical nor diagonal. In order to analyze the dynamics of fields in such a curved field-space, we resort to the conventional method for multifield inflation in a curved field-space [42, 43, 44]. From now on we work in the basis of $\phi^{a=1,2}\equiv(\phi_{r},\phi_{i})^{T}$ unless otherwise stated. We define the Levi-Civita connection, covariant field-derivative $\nabla_{a}$, and covariant time-derive in the field space as
@@ 3.1 Dynamics of the homogeneous mode
One can solve numerically these equations of motion of fields straightforwardly. However, it is instructive to see a limiting case which can be proven simpler to understand. For that, one can take a look at the case of $\alpha\lll 1$ for which approximate analytic expressions can be obtained as follows. The frame function in $\left(\phi,\theta\right)$-basis is approximated as
@@
$K_{22}$$\approx$ $\phi^{2}+\frac{\left[1+6\xi_{\phi}\left(\alpha\cos(2\theta)\right)^{2}\right]\xi_{\phi}\phi^{4}}{M_{\rm P}^{2}}\ ,$
$K_{12}$$\approx$ $6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\phi\alpha\cos\left(2\theta\right)\ll 1\ .$ (3.13)
@@
$H_{I}\simeq\frac{\lambda_{\Phi}^{1/2}M_{\rm P}}{2\sqrt{3}\xi_{\phi,0}}$ (3.20)
@@
From Eq. (3.22), the mass-squared along $\varphi_{2}$-direction is found to be
@@
$\alpha_{\rm c}\equiv 1/(24\xi_{\phi,0}),$ (3.25)
the mass along $\varphi_{2}$ direction (i.e., angular direction in the complex field space of $\Phi$) is comparable to the expansion rate given in Eq. (3.20). This implies that, in the large field region where the potential along the radial direction (i.e., along $\varphi_{1}$ or $\phi$) becomes quite flat, if $\alpha\gtrsim\alpha_{\rm c}$, the dynamics of field would be initially along $\varphi_{2}$ toward the vicinity of $\varphi_{2}\sim 0$ and afterward moves nearly along $\varphi_{1}$ toward the origin, as will be shown by numerical analysis. On the other hand, if $\alpha\lll\alpha_{\rm c}$, the trajectory will be nearly radial toward the origin. That is, the trajectory of the field would look like one of the following three cases in $(\phi_{r},\phi_{i})$ space:
@@
Figure 2: Trajectories of the field dynamics for $\alpha=10^{-4},10^{-3},10^{-2},10^{-1}$ from left to right with $\xi_{\phi,0}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/16,\pi/4,7\pi/16$(from left to right).
@@
As it can be seen from the dynamics of the inflaton with $\theta_{\rm ini}\neq 0$, a salient and interesting feature of our hybrid axi-Majoron inflation is that the PQ field $\Phi$ can have angular motion even after inflation, if $\alpha$ is neither too large nor too small. Figure 3 shows the time-dependence of $\theta$ and its time-derivative.
Figure 3: The time dependence of the phase $\theta$ (left) and its time-derivative (right) for $\alpha=10^{-3}$ with $\xi_{\phi,0}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$. $H_{\rm ref}\equiv\sqrt{U(\phi_{\rm ini},\theta_{\rm ini})}/\sqrt{3}M_{\rm P}\approx H_{I}$. The vertical dotted line indicates the end of inflation.
The angular momentum in a comoving volume is conserved after inflation, since the symmetry-breaking effect becomes negligible as $\phi$ becomes smaller than the Planck scale. This corresponds to the asymmetry of $U(1)_{\rm PQ}$-charged particle $\Phi$ as in the case of Affleck-Dine mechanism [29]  [^13] In Ref. [46], generation of Peccei-Quinn number asymmetry and subsequent baryogenesis was discussed, but in a different realization.. The asymmetry can be transferred to the visible sector, as the baryon number asymmetry, via the usual way of Affleck-Dine mechanism and/or the mechanism of spontaneous baryogenesis [47]. More details will be discussed in Sec. 4.
@@ 3.2 Perturbations
For the estimations of perturbations [^14] There are perturbations of $\chi$ field too. However, for $\xi_{\chi}\ll\xi_{r,i}$ the background dynamics is nearly along $\Phi$ with barely changing $\chi$. In addition, in a simplified two field case with negligible kinetic mixing which is the case of $\xi_{\chi}\lll 1$, one can see that the growing rate of isocurvature perturbation is negative (i.e., suppression instead of growth) in the slow-roll and slow-turn limit [48]. Hence, we can ignore the perturbations of $\chi$ in our discussion. , we take the Newtonian gauge in an isotropic universe for the metric perturbations, i.e., the perturbed metric is given by
@@
$\epsilon$$\equiv$ $-\frac{\dot{H}}{H^{2}}=\frac{\dot{\phi}_{I}^{2}}{2H^{2}M_{\rm P}^{2}}$ (3.30)
@@
¨$\ddot{\mathcal{R}}+\left(3+2\epsilon-2\eta_{T}\right)H\dot{\mathcal{R}}+\frac{k^{2}}{a^{2}}\mathcal{R}$ $=$ $-2\dot{\theta}_{T}\frac{H}{\dot{\phi}_{I}}\left[\dot{Q}_{N}+\left(3-\eta_{T}-2\epsilon+\frac{\dot{\eta}_{N}}{H\eta_{N}}\right)HQ_{N}\right]$ (3.37)
@@
For numerical analysis, we took $\xi_{\phi,0}=1$, $\alpha=10^{-3}$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$ as an illustrative set of parameters.
@@
Figure 4: Left: the turning rate, $\dot{\theta}_{T}(t)$, for $t\leq t_{e}$. Right: the effective mass-squared $m_{\rm eff}^{2}$ associated with isocurvature perturbations for the same parameter set as the one used in Fig. 3. The vertical dotted line corresponds to $52$ $e$-folds to the end of inflation at $t=t_{e}\simeq 352H_{\rm ref}^{-1}$. Roughly, it corresponds to the time when the comoving scale associated with the pivot scale of the Planck mission exits the horizon during inflation.
In Fig. 4, $\dot{\theta}_{T}$ and $m_{\rm eff}^{2}$ are depicted. From the left panel of the figure, one can see that, $|\dot{\theta}_{T}|=\mathcal{O}(10^{-3})\times H_{I}$, i.e., the turning rate of the inflaton trajectory is quite small. Also, the right panel shows that, $m_{\rm eff}^{2}/H_{I}^{2}\lesssim\mathcal{O}(10^{-2})$ when large cosmological scales of our interest exit horizon, but it increases rapidly and becomes of $\mathcal{O}(0.1-1)$ within the last about 10 $e$-folds before the end of inflation.
@@
When $\dot{\theta}/H\lll 1$ during inflation that is the case in our scenario, the source terms in Eqs. (3.37) and (3.38) (i.e., the right-hand sides of the equations) are negligible, and both $\mathcal{R}$ and $Q_{N}$ are nearly frozen soon after crossing the horizon although there appears a mild suppression of $Q_{N}$ due to the increasing $m_{\rm eff}^{2}$ toward the end of inflation. Thus, there is no sizable impact of a two-dimensional curvilinear trajectory of the inflaton on the evolution of adiabatic and isocurvature perturbations associated with cosmological scales of the present universe, and non-Gaussianities are not enhanced [53] either. Also, even though the inflaton trajectory involves axi-Majoron direction which has anomalous couplings to non-Abelian gauge fields of the SM, the perturbations of gauge fields are not enhanced because the velocity of axi-Majoron during inflation is maintained to be smaller than the expansion rate by at least a couple of orders of magnitude during the whole relevant $e$-folds, as shown in Fig. 3.
In Fig. 5, we depict the evolutions of $Q_{N},\mathcal{R}$, and $\mathcal{S}$ for comoving scales exiting the horizon at $e$-folds of $N=37$(dashed) and $52$(solid) before the end of inflation as examples.
Figure 5: Left: the evolution of $Q_{N}$(the field-perturbation orthogonal to the inflaton direction). Right: the evolution of $\mathcal{R}$)(blue) and isocurvature($\mathcal{S}$)(red) perturbations. In both panels, solid and dashed lines correspond to the modes for the comoving scale $k_{N}=a_{N}H_{I}$ associated with the $e$-folds of $N=52$(solid) and $37$(dashed) before the end of inflation. The parameter set used for the plots is the same as the one in Fig. 3. The initial condition for the numerical integration of mode equations in Eqs. (3.37) and (3.38) was set as $Q_{T}(t_{N})=Q_{N}(t_{N})=H(t_{N})/2\pi$ and ˙$\dot{\mathcal{R}}(t_{N})=\dot{Q}_{N}(t_{N})=0$ where $H(t_{N})$ is the expansion rate when a comoving scale $k_{N}$ exit the horizon at $t_{N}$.
From the figure, one can see that, while the adiabatic ones are barely affected, isocurvature perturbations are suppressed mainly because $S=Q_{N}/\sqrt{2\epsilon}M_{\rm P}\propto 1/\sqrt{2\epsilon}$. The uncorrelated isocurvature perturbations is constrained as [54]
$\beta_{\rm iso}\equiv\frac{P_{\mathcal{S}}}{P_{\mathcal{R}}+P_{\mathcal{S}}}<3.8\times 10^{-2}$ (3.41)
where $P_{\mathcal{R}}$ and $P_{\mathcal{S}}$ stand for the power spectra of adiabatic and isocurvature perturbations, respectively. Hence, the potential problem of axion isocurvature perturbation for the relevant cosmological scales of CMB observations is absent in our scenario.
@@
$P_{\mathcal{R}}$ $=$ $\frac{1}{8\pi^{2}\epsilon}\times\frac{H^{2}}{M_{\rm P}^{2}}\ ,$ (3.43)
@@
There might be a question about the impact of isocurvature perturbations on adiabatic ones as discussed in Refs. [57, 58] for examples. Including the contribution, the adiabatic perturbation may be written as
$\mathcal{R}=\mathcal{R}_{0}+\Delta\mathcal{R}\ ,$ (3.46)
where $\mathcal{R}_{0}$ is the pure adiabatic contribution, and $\Delta\mathcal{R}$ is the one from isocurvature perturbations. In our scenario, ignoring the effect of the effective mass-square ($m_{\rm eff}^{2}$) of the isocurvature modes, we may approximate $\Delta\mathcal{R}$ as
[…]
where the integration is with respect to the $e$-folding number after the horizon exit of the mode of interest to the end of inflation. For the uncorrelated isocurvature perturbations, the power spectrum $P_{\mathcal{R}}$ is then given by
$P_{\mathcal{R}}=P_{\mathcal{R}}^{0}\times\left[1+\Bigg{\langle}\left(\frac{\Delta\mathcal{R}}{\mathcal{R}_{0}}\right)^{2}\Bigg{\rangle}\right]\ ,$ (3.48)
where $P_{\mathcal{R}}^{0}$ is the pure adiabatic contribution to the power spectrum, and the contribution to the bare spectral index $n_{s}^{0}$ of the adiabatic modes only is given by
$\Delta n_{s}\equiv n_{s}-n_{s}^{0}=2\left(\frac{\Delta\mathcal{R}}{\mathcal{R}_{0}}\right)\left[\frac{1}{\mathcal{R}_{0}}\frac{d\Delta\mathcal{R}}{dN}-\left(\frac{\Delta\mathcal{R}}{\mathcal{R}_{0}}\right)\frac{d\ln\mathcal{R}_{0}}{dN}\right]\ .$ (3.49)
Ignoring the minor changes of $Q_{N}$ and $\dot{\theta}_{T}/H$ during inflation, one finds
[…]
$\frac{d\Delta\mathcal{R}}{\mathcal{R}_{0}dN}$$\approx$ $-\frac{2\dot{\theta}_{T}}{H}\lesssim\mathcal{O}(10^{-3})$ (3.51)
with $d\ln\mathcal{R}_{0}/dN\approx\left(n_{s}^{0}-1\right)/2=-\mathcal{O}(10^{-2})$. Hence, it is expected that $\Delta n_{s}=\mathcal{O}(10^{-5}-10^{-4})$ which is smaller than the $1-\sigma$ error in the result of the Planck satellite experiment [54] at least by about an order of magnitude, and we ignore the contribution in the subsequent discussion.
For the potential given in Eq. (3.19) with $\alpha\ll 1$, the relations between slow-roll parameters and $e$-folds $N$ in the SM Higgs-inflation [45] are expected still to hold at least approximately. The latter is given by [45]
$N(\phi)\approx\frac{3}{4}\frac{\phi^{2}-\phi_{e}^{2}}{M_{\rm P}^{2}/\xi_{\phi}}$ (3.52)
as the $e$-folds from the end of inflation at $\phi_{e}\approx\left(4/3\right)^{1/4}M_{\rm P}/\sqrt{\xi_{\phi,0}}$. Slow-roll parameters are then given by
$\epsilon$ $\simeq$ $\frac{3}{4}N^{-2}\ ,$ (3.53)
$\eta_{T}$ $\simeq$ $N^{-1}\ ,$ (3.54)
@@
$P_{\mathcal{R}}$ $\simeq$ $\frac{\lambda_{\phi}N^{2}}{72\pi^{2}\xi_{\phi,0}^{2}}\ ,$ (3.55)
$n_{s}$ $\simeq$ $1-2N^{-1}\ ,$ (3.56)
$r_{T}$ $\simeq$ $12N^{-2}\ .$ (3.57)
Density perturbations of the universe extracted from various observations including CMB observations lead to $P_{R}\simeq 2.1(\pm 0.03)\times 10^{-9}\ (1\sigma\ {\rm CL})$, $n_{s}=0.9658\pm 0.004\ (1\sigma\ {\rm CL})$, and $r_{T}<0.068\ (2\sigma\ {\rm CL})$ for the pivot scale, $k_{*}=0.05/{\rm Mpc}$ [54]. Hence, Eq. (3.55) constrains $\lambda_{\phi}$ and $\xi_{\phi,0}$ to satisfy a relation,
$\lambda_{\phi}\sim 6\times 10^{-10}\xi_{\phi,0}^{2}\left(\frac{52}{N_{*}}\right)^{2}\ .$ (3.58)
where we used $N_{*}=52$ as a representative value of $e$-folds associated with $k_{*}$. In Table 2, we show a result of numerical estimations for some sets of parameters set as
$\lambda_{\phi}\equiv\lambda_{\phi}^{\rm ref}\times\xi_{\phi,0}^{2},\ \alpha=\frac{10^{-3}}{\xi_{\phi,0}},\ \phi_{\rm ini}=\frac{20M_{\rm P}}{\sqrt{\xi_{\phi,0}}},\ \theta_{\rm ini}=\frac{\pi}{4}\ ,$ (3.59)
with $\lambda_{\phi}^{\rm ref}\equiv 6\times 10^{-10}$ and several different values of $\xi_{\phi,0}$ but $N_{*}=52$ fixed.
$\xi_{\phi,0}$ $P_{R}\times 10^{9}$ $n_{s}$ $r_{T}\times 10^{3}$
@@
The result is in a good agreement with approximations in Eq. (3.55) – (3.57), and can match well with observations by adjusting $\lambda_{\phi}$. The sizable difference between the cases of $\xi_{\phi,0}=1$ and the other much larger values may be understood from the dependence of $U$ on $\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ in Eq. (3.19). For convenience, in the subsequent argument we use $\lambda_{\phi}$ parametrized as in Eq. (3.59).
@@
$\xi_{\phi,0},\lambda_{\phi},\alpha,\theta_{\rm ini},\phi_{\rm ini}\ .$ (3.60)
Once $\xi_{\phi,0}$ as a free parameter is chosen such that $\xi_{\phi,0}\gtrsim\mathcal{O}(1)$, $\lambda_{\phi}$ is nearly fixed as long as $\alpha$ is sufficiently smaller than unity. Dependence of observables on the other parameters for given $\xi_{\phi,0}$ and $\lambda_{\phi}$ is shown in Fig. 6.
Figure 6: $P_{R}$(left) and $n_{s}$(right) as functions of $\theta_{\rm ini}$ for various values of $\alpha$ with $\xi_{\phi,0}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}/M_{\rm P}=10({\rm top}),20({\rm bottom})$.
It turns out that inflationary observables are barely affected by $\theta_{\rm ini}$ for $\sqrt{\xi_{\phi,0}}\phi_{\rm ini}\gtrsim 10M_{\rm P}$ as long as $\alpha\xi_{\phi,0}\lesssim\mathcal{O}(10^{-2})$ except the case of $\theta_{\rm ini}$ close to $\pi/2$ with $\sqrt{\xi_{\phi,0}}\phi_{\rm ini}\sim 10M_{\rm P}$ when $\alpha\xi_{\phi,0}\sim 10^{-2}$. Note that, if angular dynamics is not strong, for a given $e$-folds associated with the pivot scale of observations, the spectral index is nearly fixed irrespective of initial position of the field dynamics. The amplitude of power spectrum has only minor dependence on the initial position, and the dependence can be compensated by adjusting $\lambda_{\phi}$ without affecting the spectral index. This is because, while the former is a function of $e$-folds only, the latter is a function of $e$-folds and energy density of inflation. On the other hand, if $\alpha\xi_{\phi,0}\gtrsim\mathcal{O}(10^{-1})$, the observables varies a lot, depending on $\theta_{\rm ini}$, but $\phi_{\rm ini}$-dependence becomes weak if it is sufficiently larger than the minimal value to accommodate about 60 $e$-folds.